OpenStudy (anonymous):

Solve for x: 4/5 x + 4/3 = 2x Solve for x: 3 (x+1) = -2 (x-1) - 4

OpenStudy (goformit100):

Use Elimination Method.

OpenStudy (anonymous):

for the second one: 3(x+1)=-2(x-1)-4 3x+3=-2x+2-4 3x+3=-2x-2 3x=-2x-5 5x=-5 5x/5=-5/5 x=-1

OpenStudy (anonymous):

Thank you! @juliax

OpenStudy (anonymous):

you're welcome :) im trying to work out the first one ;3

OpenStudy (anonymous):

i came to an answer but im honestly not sure it's right, i got \[\sqrt{32/30}=x \]

OpenStudy (anonymous):

can you help me with a problem when your done satellite?

OpenStudy (anonymous):

\[\frac{4}{5}x+\frac{4}{3}=2x\] multiply both sides by 15 and get \[12x+16=30x\] then solve for \(x\)

OpenStudy (phi):

sat has the right idea, but his arithmetic is rusty. do what he said, multiply both sides by 15, but notice (4/3) * 15 is not 16

OpenStudy (anonymous):

phi can you help me with a question? Evaluate the following expression using the given values: Find x - 3y if x = 3 and y = -2. -9 -3 3 9